OpenStudy (anonymous):

Solve √ 2x+3=x+1 for x.

OpenStudy (anonymous):

X^2-2

OpenStudy (mimi_x3):

is it: \(\sqrt{2x+3}=x+1 \)

OpenStudy (anonymous):

√ 2x+3=x+1 √ 2x=x+1-3 2x = (x-2)^2 x = (x-2)^2 / 2

OpenStudy (anonymous):

Thanks you guys!

OpenStudy (anonymous):

:)